OpenStudy (anonymous):

Dan used a probability simulator to roll a 6-sided number cube and flip a coin 100 times. The results are shown in the tables below: Number on the Cube Number of Times Rolled 1 19 2 26 3 13 4 23 5 8 6 11 Heads Tails 48 52 Using Dan's simulation, what is the probability of rolling a 4 on the number cube and the coin landing heads up?

OpenStudy (kropot72):

The experimental probability of rolling a four is 23/100. The experimental probability of heads is 48/100. The events 'roll a four' and 'heads up' are independent, therefore \[\large P(4 \cap heads)=\frac{23}{100}\times\frac{48}{100}\]